OpenStudy (anonymous):

Find g.s \[\cos 3x -\cos x + \sin 4x=0\]?

OpenStudy (anonymous):

just continue

OpenStudy (anonymous):

-2 sin2x sinx + sin 4x =0

OpenStudy (anonymous):

then??

OpenStudy (anonymous):

Ouch, sorry don't know

OpenStudy (anonymous):

\[-2\ \sin 2x \sin x+2 \sin 2x \cos 2x=0\]\[\sin 2x(\cos 2x-\sin x)=0\]

OpenStudy (anonymous):

\[\sin 2x=0\]or\[\cos 2x-\sin x=0 \rightarrow 2\sin^2 x+\sin x-1=0\] im watching olympic

hartnn (hartnn):

so sin x is 0.5 or x is 30....overall,x=0,30 degrees...

OpenStudy (anonymous):

hey isnt this my that i posted earlier? :/

OpenStudy (anonymous):

my question*

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

i also have doubts so i posted

OpenStudy (anonymous):

sinx (2sinx +1) = sin 90

OpenStudy (anonymous):

if\[\sin 2x=0\]\[2x=n\pi\] so \[x=\frac{n\pi}{2}\] ----------------------------------- if \[2 \sin^2 x+\sin x-1=0\] make a quadratic in terms of \(\sin x\) and enjoy

OpenStudy (anonymous):

but your working looks nicer than mine..my one got ugly :/

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

wait wait i think you forgot about the cos3x

OpenStudy (anonymous):

sin x = -1 or sinx = 1/2

OpenStudy (anonymous):

@mukushla is that correct

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thxx

OpenStudy (anonymous):

np...:)